A ball is projected vertically upward from the top of a tower 60m high with a velocity if 30ms.what is the maximum height above the ground level? How long does it take to reach the ground

1
Expert's answer
2021-11-09T17:53:26-0500

"h=v^2/(2g)=30^2/(2\cdot9.8)=45.9\ (m)"


"H=60+45.9=105.9\ (m)"


"t=v/g+\sqrt{2h/g}=30/9.8+\sqrt{2\cdot105.9/9.8}=7.71\ (s)"